# Decoder Performance in Hybrid CV-Discrete Surface-Code Threshold Estimation Using LiDMaS+

### 06 Mar 2026 v1

@@ I.3 LiDMaS+ as the Experimental Platform
Reproducibility is supported through deterministic seeding, fixed sweep grids, and structured outputs. Each run records the chosen seed, the decoder, the distance, the physical-noise parameter, the number of trials, and the resulting logical error statistics in machine-readable tabular form. The reported outputs also include confidence intervals and auxiliary decoder diagnostics such as defect and weight summaries, which makes it possible to reproduce both the figures and the threshold summary tables from the saved artifacts rather than from hand-copied numbers. For threshold-oriented Pauli runs, LiDMaS+ further emits scaling reports and JSON summaries used to construct the finite-size table. In the hybrid case, the same tabular outputs support a more conservative crossing analysis derived directly from the sampled curves.
The specific execution path used in the paper is the dedicated paper_runs/ workflow added on top of the repository examples. That layer does not alter the decoding logic. It standardizes the experiment matrix, output directories, merged tabular-data generation, and publication-figure rendering. This separation is useful because it keeps the scientific workflow explicit and auditable while preserving LiDMaS+ itself as the common simulation backend. In that sense, LiDMaS+ is suitable for the present study not because it automates every possible analysis, but because it provides a stable and decoder-consistent environment in which methodological differences can be isolated and compared.
@@ II.3 Threshold Estimation Protocol
All experiments follow the same basic protocol. For a chosen decoder, code distance, and point on the physical-noise sweep, LiDMaS+ generates repeated noisy samples, performs syndrome extraction and decoding, and records whether each trial results in a logical failure. The logical error rate is then estimated as the empirical failure frequency at that operating point, and the simulator reports accompanying confidence intervals in the output tables. This procedure is repeated across a matched sweep grid for every decoder so that each curve comparison is pointwise aligned in both noise parameter and trial count.
@@ II.5 Metrics
The primary metric throughout the paper is the logical error rate (LER). This is the quantity plotted in the fixed-distance and multi-distance figures, and it is the metric from which all threshold-oriented summaries are derived. Confidence intervals are used to judge whether observed curve separations are likely to reflect a real decoder effect rather than pure sampling noise. For the Pauli threshold study, we additionally report crossing estimates, collapse-based critical-point estimates, the fitted exponent $\nu$, and the collapse cost returned by the scaling routine. For the hybrid threshold study, we report decoder-labeled crossing proxies derived from the sampled multi-distance curves, while explicitly noting when those summaries are limited by grid resolution.
@@ III Experimental Design
This design leads directly to the figure and table structure used in the paper. The fixed-distance Pauli and hybrid runs supply the baseline comparison figure. The hybrid multi-distance run supplies the main scaling-style figure together with a decoder-labeled summary table. The Pauli threshold run supplies the finite-size summary table, while the hybrid threshold run supplies a practical crossing-summary table derived from the sampled data.
@@ IV Results
We now report the decoder-comparison results obtained from the fixed-distance baseline experiments, the multi-distance hybrid sweeps, and the threshold-oriented Pauli and hybrid analyses. All fixed-distance and multi-distance curve comparisons were generated with deterministic seeds and matched sweep grids. The Pauli baseline figure uses $3000$ trials per point, while the hybrid baseline, hybrid multi-distance, and threshold-oriented runs use $2000$ trials per point (with $100$ bootstrap samples for the Pauli scaling analysis). The main empirical message is that decoder choice clearly matters in the Pauli baseline and remains visible in the hybrid setting: Union-Find degrades relative to MWPM, while trained neural-guided MWPM stays close to MWPM at moderate noise but shows elevated decoder-failure diagnostics at the highest-noise points.
@@ IV.4 Comparison with Previous Studies
To position the present decoder-comparison results in a broader context, Table 6 summarizes representative published values from APS, Nature, and Springer venues together with the corresponding LiDMaS+ outcomes. Because these studies differ in noise model (Pauli, biased-noise, and hardware-realistic), decoder assumptions, and endpoint metrics, the values should be read as contextual anchors rather than as strict one-to-one rankings.

### 25 Mar 2026 v2

@@ I.3 LiDMaS+ as the Experimental Platform
Reproducibility is supported through deterministic seeding, fixed sweep grids, and structured outputs. Each run records the chosen seed, the decoder, the distance, the physical-noise parameter, the number of trials, and the resulting logical error statistics in machine-readable tabular form. The reported outputs also include confidence intervals and auxiliary decoder diagnostics such as defect and weight summaries, enabling full reconstruction of the reported threshold summaries and curve-level conclusions directly from saved artifacts rather than hand-copied numbers. For threshold-oriented Pauli runs, LiDMaS+ further emits scaling reports and JSON summaries for finite-size analysis. In the hybrid case, the same tabular outputs support a more conservative crossing analysis derived directly from the sampled curves.
The specific execution path used in the paper is the dedicated examples/paper_runs/paper_01 workflow added on top of the repository examples. That layer does not alter the decoding logic. It standardizes the experiment matrix, output directories, and merged tabular-data generation for downstream analysis. This separation is useful because it keeps the scientific workflow explicit and auditable while preserving LiDMaS+ itself as the common simulation backend. In that sense, LiDMaS+ is suitable for the present study not because it automates every possible analysis, but because it provides a stable and decoder-consistent environment in which methodological differences can be isolated and compared.
@@ II.3 Threshold Estimation Protocol
All experiments follow the same basic protocol. For a chosen decoder, code distance, and point on the physical-noise sweep, LiDMaS+ generates repeated noisy samples, performs syndrome extraction and decoding, and records whether each trial results in a logical failure. The logical error rate is then estimated as the empirical failure frequency at that operating point, and the simulator reports accompanying confidence intervals in structured outputs. This procedure is repeated across a matched sweep grid for every decoder so that each curve comparison is pointwise aligned in both noise parameter and trial count.
@@ II.5 Metrics
The primary metric throughout the paper is the logical error rate (LER). This metric is tracked across both fixed-distance and multi-distance analyses, and all threshold-oriented summaries are derived from it. Confidence intervals are used to judge whether observed curve separations are likely to reflect a real decoder effect rather than pure sampling noise. For the Pauli threshold study, we additionally report crossing estimates, collapse-based critical-point estimates, the fitted exponent $\nu$, and the collapse cost returned by the scaling routine. For the hybrid threshold study, we report decoder-labeled crossing proxies derived from the sampled multi-distance curves, while explicitly noting when those summaries are limited by grid resolution.
@@ III Experimental Design
This design maps each run block to a distinct analysis outcome. The fixed-distance Pauli and hybrid runs provide baseline decoder comparisons. The hybrid multi-distance run provides the distance-scaling comparison with decoder-labeled summaries. The Pauli threshold run provides finite-size threshold summaries, while the hybrid threshold run provides a practical crossing summary derived from sampled data.
@@ IV Results
We now report the decoder-comparison results obtained from the fixed-distance baseline experiments, the multi-distance hybrid sweeps, and the threshold-oriented Pauli and hybrid analyses. All fixed-distance and multi-distance curve comparisons were generated with deterministic seeds and matched sweep grids. The Pauli baseline run uses $3000$ trials per point, while the hybrid baseline, hybrid multi-distance, and threshold-oriented runs use $2000$ trials per point (with $100$ bootstrap samples for the Pauli scaling analysis). The main empirical message is that decoder choice clearly matters in the Pauli baseline and remains visible in the hybrid setting: Union-Find degrades relative to MWPM, while trained neural-guided MWPM stays close to MWPM at moderate noise but shows elevated decoder-failure diagnostics at the highest-noise points.
@@ IV.4 Comparison with Previous Studies
To position the present decoder-comparison results in a broader context, Table 6 summarizes representative published values from literature with the corresponding LiDMaS+ outcomes. Because these studies differ in noise model (Pauli, biased-noise, and hardware-realistic), decoder assumptions, and endpoint metrics, the values should be read as contextual anchors rather than as strict one-to-one rankings.